An alkane is optically active. What should be its minimum molar mass?

Grade:12

2 Answers

lokesh palingi
10 years ago
Alkanes belong to a homologous series of organic compounds in which the members differ by a molecular mass of 14.03U (mass of a methanediyl group, —CH2—, one carbon atom of mass 12.01u, and two hydrogen atoms of mass ˜1.01u each),there fore molecular mass of Alkanes is (12.01U+2*(1.01)=12.01+2.02=14.03U)
Tejasv Gaur
6 years ago
To be optically active an Alane should have atleast 7 carbons i.e. C7H16. So, molar mass » 7*12 + 16*1=100.